OpenStudy (anonymous):

Frances has 4 square baking dishes. Each pan is 7 inches × 7 inches × 2 inches. Which expression will give her the total volume of all the pans?

OpenStudy (pfenn1):

Volume of one pan is (x^2)h or 7*7*2. Volume of all pans = 4(x^2)h

OpenStudy (kropot72):

V = 4 * (7*7*7) cubic inches

OpenStudy (pfenn1):

@kropot72, I think you meant V = 4 * (7*7*2)cubic inches

OpenStudy (anonymous):

Lol cause that kinda confused me

OpenStudy (kropot72):

Sorry pfenn1. You are right. I should have been more careful :(

OpenStudy (pfenn1):

It was just a typo.